# Two-stage Design for Failure Probability Estimation with Gaussian Process Surrogates

### v2

@@ 2.2 Contour location
To facilitate failure probability estimation (4), a surrogate must accurately classify passes and failures. Strategic sequential designs targeting the failure contour, i.e., “contour location,” will outperform space-filling counterparts in this regard. In CL, after initializing with a small space-filling design, subsequent training locations are selected iteratively through the optimization of an acquisition criterion. The computer simulation is evaluated at the selected input(s), the surrogate is updated, and the process is repeated until the budget is exhausted or a stopping criterion is met.
@@
where $p_{\mathbf{x}}$ represents the probability of failure at input $\mathbf{x}$. With a GP surrogate, this probability boils down to a Gaussian CDF computation:
$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
Returning to the Herbie example, the entropy surface for the trained GP is shown in the right panel of Figure 1 and along the slice in Figure 2 (green dotted/dashed). Notice the high entropy regions mostly agree with the locations of the true failure contour (we could improve upon this with a CL design instead of an LHS). A reasonable acquisition could fall in any of these local optima.
Naturally, implementations of entropy as an aqcuisition function vary on the finer details. For GP surrogates we will leverage the optimization scheme of Cole et al., (2023), which first deploys a small LHS, identifies the LHS sample with the highest entropy, then runs a numerical optimization of entropy initialized at that point. By avoiding multi-starts, Cole et al., target local optima in the entropy surface. For DGP surrogates, we will follow Booth et al., (2024b). To avoid cumbersome numerical optimizations with the heftier MCMC-based DGP, Booth et al., use triangulation candidates (Gramacy et al.,, 2022), then select the candidate on the Pareto front of entropy and uncertainty. Henceforth, all GPs and DGPs we consider will be trained with these two contour location methods, respectively.
Figure 3: Various designs for estimating the failure probability of the Herbie function with a GP surrogate and a total budget of $N=150$. All methods start with the same CL design of size $n=100$ (black circles), leaving a remaining budget of $B=50$ (blue triangles) for either continued surrogate training or importance sampling. Left: Entire budget is spent on continued contour location, with acquisitions exploring the entire space. Center: Remaining budget is sampled from a bias distribution for importance sampling (black circles omitted since they do not feature in the SIIS estimate). Right: Remaining budget is used for MC samples with highest entropy, solely exploiting near the predicted contour. Comparison of failure probability estimates obtained from these is provided later in Section 4. Although there is no clear distinction in performace between the left/right designs for this simple illustrative example, the continued exploration of the “All CL” design will hinder performance on larger, harder problems (Figure 6).
Continuing the Herbie function example, the left panel of Figure 3 shows the design points chosen from entropy-based contour location with a GP surrogate (Cole et al.,, 2023), indicated by both the black circles and blue triangles (the distinction between the two will become clear when we discuss the other panels). Notice, the chosen points attempt to balance exploration (points spread out over the entire domain) and exploitation (points hovering around the red failure contour). The resulting surrogate may inform an $\hat{\alpha}_{\textrm{SURR}}$ estimate (4); we will include this as a competitor in our later exercises.
@@ 2.3 Importance sampling
Monte Carlo approximation of this integral provides the estimator
@@
The “bias weights” $w_{i}$ account for the difference between the input density and the bias density. This estimator is theoretically unbiased and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002). But estimation of an effective bias distribution is tricky. The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but neither $\alpha$ nor the indicator function are known a priori. For the sake of feasibility, $q(\mathbf{x})$ is traditionally constrained to a known distributional family, such as a Gaussian mixture model (GMM; Reynolds et al.,, 2009). Bias distribution training is then relegated to learning distributional parameters (in a GMM, these include the number of mixtures, mixture weights, and component means and variances), often accomplished through cross validation or information criteria maximization.
@@
In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large Monte Carlo sample of the input distribution, then train a GMM bias distribution to these predicted failures. If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$. With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$, evaluated through the expensive simulator, and deployed in Eq. (6) to provide an estimate of the failure probability.
To illustrate, consider our running Herbie function example. Using the GP surrogate trained on the 100-point contour locating design from Figure 3 (black circles) and a large MC sample from the input distribution, we trained a GMM bias distribution to the MC samples which the surrogate predicted to be failures. The blue triangles in the center panel show 50 points sampled from this bias distribution. These samples may be evaluated through the true model to estimate $\hat{\alpha}_{\textrm{IS}}$ (6), thereby exhausting our entire budget of 150 evaluations. While the blue triangles target the failure region, they do not provide enough coverage to ensure a precise estimate. In our later experiments (Section 4), SIIS performance is highly variable.
@@
While SIIS has been widely embraced for its unbiasedness, it has several key weaknesses. First, the bias distribution, which features heavily in the final estimate of Eq. (6), is a completely auxilary quantity. Hypothetically, two different bias distributions could produce the same samples $\left\{\mathbf{x}_{i}\right\}_{i=1}^{B}$, thus producing the same $\left\{f(\mathbf{x}_{i})\right\}_{i=1}^{B}$. But in the calculation of Eq. (6), these two bias distributions would contribute different weights, resulting in different final answers even though the training data from the computer simulator was exactly the same. O’Hagan, (1987) first pointed out this conundrum as a violation of the “Likelihood principle.” A second weakness is the restriction of $q(\mathbf{x})$ to distributional families like GMMs. Failure regions may be complex and may not fit nicely under these target distributions. Consequently, SIIS typically requires large $B$ in order to adequately cover the failure region. For context, Peherstorfer et al., (2016) entertained $B\in\left\{100,1000,10000,100000\right\}$ and found performance continued to improve as $B$ increased. Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation. [Throughout, we will borrow from our proposed method to inform budget allocation for the SIIS competitor, which is how we chose to separate the budget of Figure 3 into 100 for training and 50 for estimation.] Finally, the SIIS method does not leverage the full potential of the surrogate model. Many expensive simulator evaluations go into training the surrogate and, in the end, the surrogate is only used to identify predicted failures, with no mention of its UQ. [Notice, as portrayed in the center panel of Figure 3, the black circles are abandoned after the bias distribution is trained.]
@@ 3 Two-stage design for failure probability estimation
All in all, we are unsatisfied with the two state-of-the-art surrogate-based methods for estimating $\alpha$ which we just described: surrogate MC and SIIS. While the drawbacks to SIIS are apparent, the limitations of surrogate MC deserve more attention. To investigate further, we conducted the same CL sequential design of the Herbie function (Figure 3), this time storing the surrogate MC estimate $\hat{\alpha}_{\textrm{SURR}}$ (4), which we now denote as simply $\hat{\alpha}$, after every 10th acquisition. Progress is shown in the left panel of Figure 4. While there is significant learning occurring with early acquisitions, it quickly plateaus. By the time the surrogate has been trained on 100 points, it knows just as much about $\alpha$ as it does when it has been trained on the full budget of $N=150$ points. In this case, there is bound to be a more effective way to use the remaining simulation budget. We propose a solution: first developing a stopping criterion to decide when to halt contour location, then proposing a hybrid Monte Carlo estimator that greedily leverages the remaining simulation budget precisely for failure probability estimation.
Figure 4: Left: Surrogate MC failure probability estimate for the Herbie function, updated after every 10th contour locating acquisition. True $\alpha$ in dashed red. Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green. Blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
@@ 3.1 A stopping criterion for contour location
Let $n$ denote the budget used on contour location and $B$ denote the remaining budget such that the total budget is $N=n+B$. We seek an objective, automatable decision rule for selecting $n$. The right panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$ estimate from the left panel. As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero. How small of an update then is small enough to indicate learning has plateaued?
The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$, the true failure probability, and $M$, the size of the MC sample used. Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$. To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update. Notice in the estimator of Eq. (2), the indicator $\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$. Thus, the MC estimator has variance inversely proportional to $M$:
@@
Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator. This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$. Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate.
The green dashed lines in the right panel of Figure 4 show ±$\pm\hat{\sigma}_{\alpha}$, updated along with $\hat{\alpha}$. At the start, updates to $\hat{\alpha}$ are far beyond one standard error. But after a while, the updates sit comfortably within these bounds. Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$. We posit that, for fixed $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we propose halting CL once we have observed two successive updates within ±$\pm\hat{\sigma}$. The blue square in Figure 4 marks this point, stopping CL for this example at $n=100$, effectively identifying a point where surrogate learning has saturated and leaving a remaining budget of $B=50$ (this informed the budget allocation of Figure 3).
Note, this stopping criterion requires the same $X_{M}$ be used in every $\hat{\alpha}$ calculation. If a new $X_{M}$ is generated for each estimate, there will be additional variability among the $\hat{\alpha}$’s, and the proposed ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased. Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
Our budgeting scheme is summarized as follows. Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$. At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as seen in Figure 4). To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted. If the contour location design is succeeding, it will be placing some points in the failure region. We recommend a default of 10 failures, but suggest increasing this for more complex, higher dimensional problems. Once the design has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition. Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$. When this condition has been satisfied twice in a row, halt the sequential design and proceed to the next stage (the topic of Section 3.2).
@@
In practice, it is important to select a MC sample size that is large enough without being too cumbersome. Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq. 3) for every sample in $X_{M}$. Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for extremely large $M$, particularly with a heftier surrogate like a Bayesian DGP. If we knew $\alpha$, we could determine $M$ by thresholding the standard error $\sigma_{\alpha}$. For example, to achieve a standard error less than 10% of $\alpha$ requires:
@@
3.2 Hybrid Monte Carlo for failure probability estimation
Once $n$ is determined, how should we spend the remaining budget of simulator evaluations ($B=N-n$), in conjunction with our trained surrogate, to most effectively estimate $\alpha$? We propose an upgrade to the surrogate MC estimator of Eq. (4). Let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive). Then, Eq. (1) may be partitioned as

### v3

@@ 2.2 Contour location
To facilitate failure probability estimation (4), a surrogate must accurately classify passes and failures. Strategic sequential designs targeting the failure contour, i.e., “contour location,” will outperform space-filling counterparts in this regard. In CL, after initializing with a small space-filling design, subsequent training locations are selected iteratively through the optimization of an acquisition criterion. The computer simulation is evaluated at the selected input(s), the surrogate is updated, and the process is repeated until the budget is exhausted or a stopping criterion is met. Stage 1 of Figure 1 visualizes this iterative loop. While acquisitions may be made in batches, we only consider single acquisitions here.
@@
where $p_{\mathbf{x}}$ represents the predicted probability of failure at input $\mathbf{x}$, i.e., $\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)$. With a GP surrogate, this predicted probability boils down to a Gaussian CDF computation,
$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right),$
with $\mu_{n}(\mathbf{x})$ and $\sigma_{n}(\mathbf{x})$ following Eq. (3). Revisiting the Herbie example of Section 2.1, we offer two visuals of the entropy surface for the trained GP. First, the right panel of Figure 2 shows entropy over the entire 2d space. Notice high entropy regions are predominantly near the failure contour and/or away from training data locations. Second, Figure 3 shows the entropy surface (green) along the slice $x_{2}=1.05$. Entropy is high where the credible interval (blue shading) captures the failure threshold (dotted red). It is influenced by how much of the credible interval is split above/below $t$, regardless of the interval’s width. For the highest $x_{1}$ values, surrogate uncertainty is high but entropy is low because the predicted values are far from the failure threshold. As an acquisition criterion, entropy effectively identifies regions where the surrogate is uncertain of its pass/fail prediction. A reasonable acquisition could fall in any of the local optima of the entropy surface.
Naturally, implementations of entropy as an acquisition function vary on the finer details. For GP surrogates we will leverage the optimization scheme of Cole et al., (2023), which first deploys a small LHS, identifies the LHS sample with the highest entropy, then runs a numerical optimization of entropy initialized at that point. By avoiding multi-starts, Cole et al., target local optima in the entropy surface. For DGP surrogates, we will follow Booth et al., (2024c). To avoid cumbersome numerical optimizations with the heftier MCMC-based DGP, Booth et al., use triangulation candidates (Gramacy et al.,, 2022), then select the candidate on the Pareto front of entropy and uncertainty. Henceforth, all GPs and DGPs we consider will be trained with these two contour location methods, respectively.
The right and center panels of Figure 2 showed a contour locating design for the Herbie function. Starting with a GP surrogate trained on an LHS design of size $n_{0}=20$ (white circles), we used the entropy-based scheme of Cole et al., (2023) to acquire 30 more inputs, one-at-a-time (white triangles). The acquired points attempt to balance exploring the space and exploiting learning of the failure contour.
@@ 2.3 Importance sampling
Monte Carlo approximation of this integral provides the estimator,
@@
This estimator requires $B$-many evaluations of the expensive simulator $f$. The “bias weights” $w_{i}$ account for the difference between the input density and the bias density. This estimator is theoretically unbiased for any $B>0$ and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002). But estimation of an effective bias distribution is tricky. The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but neither $\alpha$ nor the indicator function are known a priori. For the sake of feasibility, $q(\mathbf{x})$ is traditionally constrained to a known distributional family, such as a Gaussian mixture model (GMM; Reynolds et al.,, 2009). Bias distribution training is then relegated to learning distributional parameters (in a GMM, these include the number of mixtures, mixture weights, and component means and variances), often accomplished through cross validation or information criteria maximization.
@@
In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large Monte Carlo sample of the input distribution, then train a GMM bias distribution to these predicted failures. Recent works have leveraged contour location to train surrogates for this very purpose (e.g., Cole et al.,, 2023; Booth et al.,, 2024b). If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$. With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$, evaluated through the expensive simulator, and deployed in Eq. (6) to provide an estimate of the failure probability. SIIS is a two-stage enterprise (first training a surrogate then using the surrogate to estimate a bias distribution for importance sampling), but it is not a “two-stage design” since the final estimator makes no use of the surrogate nor the observations that had been used to train the surrogate.
@@
Although SIIS has been widely embraced for its unbiasedness, it has several key weaknesses. First, the bias distribution, which features heavily in the final estimate of Eq. (6), is a completely auxiliary quantity. Hypothetically, two different bias distributions could produce the same samples $\left\{\mathbf{x}_{i}\right\}_{i=1}^{B}$, thus producing the same $\left\{f(\mathbf{x}_{i})\right\}_{i=1}^{B}$. But in the calculation of Eq. (6), these two bias distributions would contribute different weights, resulting in different final answers even though the training data from the computer simulator was exactly the same. O’Hagan, (1987) first pointed out this conundrum as a violation of the “Likelihood principle.” A second weakness is the restriction of $q(\mathbf{x})$ to distributional families like GMMs. Failure regions may be complex and may not fit nicely under these target distributions. Consequently, SIIS typically requires large $B$ in order to adequately cover the failure region. For context, Peherstorfer et al., (2016) entertained $B\in\left\{100,1000,10000,100000\right\}$ and found performance continued to improve as $B$ increased. Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation. [In our benchmark exercises, we will borrow from our proposed method to inform budget allocation for the SIIS competitor.] Finally, the SIIS method does not leverage the full potential of the surrogate model. Many expensive simulator evaluations go into training the surrogate, but the surrogate is abandoned after it has identified potential bias locations for estimation of $q(\mathbf{x})$.
@@ 3 Two-stage design for failure probability estimation
Ultimately, we are unsatisfied with the two aforementioned surrogate-based methods for estimating $\alpha$: exhaustive CL with surrogate MC and surrogate-informed importance sampling. While the drawbacks to SIIS are apparent, the limitations of exhaustive CL deserve more attention. To investigate further, we revisit the CL design of the Herbie function (Figure 2), this time continuing contour location until an entire budget of $N=150$ has been exhausted. As the design progresses, we evaluate and store the surrogate MC estimate $\hat{\alpha}_{\textrm{SURR}}$ (4), which we now denote as simply $\hat{\alpha}$, after every 10th CL acquisition. Progress in this estimate is shown in the left panel of Figure 4. While there is significant learning occurring with early acquisitions, it quickly plateaus. By the time the surrogate has been trained on 100 points, it knows just as much about $\alpha$ as it does when it has been trained on the full budget of $N=150$ points. In this case, continued contour location is effectively wasting the remaining simulation budget. We propose a solution: first developing a stopping criterion to decide when to halt contour location (Figure 1, Stage 1), then proposing a second design stage (Figure 1, Stage 2) which greedily leverages the remaining simulation budget for Monte Carlo samples with the highest classification uncertainty. The second stage, in which observations are acquired directly from $p(\mathbf{x})$ enables a“hybrid Monte Carlo” estimator, using both true simulator observations and surrogate predictions.
Figure 4: Left: Failure probability estimate (4) from a GP surrogate of the Herbie function (as shown in Figure 2), updated after every 10th acquisition of the CL design. True $\alpha$ in dashed red. Right: Difference in successive $\hat{\alpha}$ estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green. Narrow teal line indicates true $\sigma_{\alpha}$. Blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$, which satisfies our proposed stopping criterion.
@@ 3.1 A stopping criterion for contour location
Let $n$ denote the budget of expensive simulator evaluations used on contour location (Stage 1). Let $B$ denote the remaining evaluation budget, such that the total budget is $N=n+B$. We seek an objective, automatable decision rule for selecting $n$, halting CL once surrogate learning has plateaued. Consider the progress in $\hat{\alpha}$ shown in Figure 4. The right panel shows the relative difference between each successive $\hat{\alpha}$ estimate from the left panel. As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but with this finite budget they never settle at exactly zero. We need an objective method to decide when these updates to $\hat{\alpha}$ are no longer significant.
The challenge is that $\hat{\alpha}$ exhibits variability from two distinct sources: surrogate error and the natural variability that accompanies any Monte Carlo estimator with finite $M$. Hypothetically, presume we could evaluate the true $f$ an unlimited number of times to obtain the brute force MC estimator of Eq. (2). In this estimator, the indicator $\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$. Thus, the MC estimator has variance inversely proportional to $M$:
@@
Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator; it quantifies the degree of variability that we would expect even if we used the true $f$ (i.e., zero surrogate error). We consider successive updates to $\hat{\alpha}$ within this realm of intrinsic MC variability to indicate stabilization of the surrogate. Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate. While this does intertwine surrogate error and MC error, it still provides a useful and feasible estimate.
To demonstrate, the green dashed lines in the right panel of Figure 4 show ±$\pm\hat{\sigma}_{\alpha}$, updated with $\hat{\alpha}$ after every 10th acquisition. The estimated $\hat{\sigma}_{\alpha}$ (dashed green) effectively reflects the true $\sigma_{\alpha}$ (narrow solid teal) after very few acquisitions. At the start of the design, updates to $\hat{\alpha}$ are far beyond this estimated standard error. But after a while, the updates sit comfortably within these bounds.
Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$. We posit that, for fixed $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is insignificant, as it is within the realm of expected MC error. [To avoid excessive computation, we recommend updating estimates after every 10 acquisitions, which we will do in all later exercises.] Accordingly, we propose halting CL once we have observed two successive updates within ±$\pm\hat{\sigma}$. The blue square in Figure 4 marks this point, stopping CL for this example at $n=100$, effectively identifying a point where surrogate learning has saturated and leaving a remaining budget of $B=50$. Note, this stopping criterion requires the same $X_{M}$ be used in every $\hat{\alpha}$ calculation. If a new $X_{M}$ is generated for each estimate, there would be additional sampling variability reflected among the $\hat{\alpha}$’s (on top of the surrogate error and MC variability).
Our budgeting scheme is summarized as follows. Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$. At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as seen in Figure 4). To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before contour location could be halted. If the contour locating design is succeeding, it will be placing some points in the failure region. We recommend a default of 10 failures, but suggest increasing this for more complex, higher dimensional problems. Once the design has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition. Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$. When this condition has been satisfied twice in a row, halt the CL design of Stage 1 and proceed to Stage 2 (the topic of Section 3.2).
@@
In practice, it is important to select a MC sample size that is large enough without being too cumbersome. Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq. 3) for every sample in $X_{M}$. Although this computation is significantly cheaper than evaluation of the expensive simulator, it could be significant for extremely large $M$, particularly with a heftier surrogate like a Bayesian DGP.
If we knew $\alpha$, we could determine $M$ by thresholding the standard error $\sigma_{\alpha}$. For example, to achieve a standard error less than 10% of $\alpha$ requires:
@@
3.2 Stage 2: Hybrid Monte Carlo for failure probability estimation
Once $n$ is determined and the contour location of Stage 1 is complete, how should we spend the remaining budget of simulator evaluations ($B=N-n$) in Stage 2? Recall, our objective is to leverage the trained surrogate to estimate $\alpha$ (1). We propose a new design for Stage 2, which enables an upgrade to the surrogate MC estimator of Eq. (4).
To set the stage, let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive). Then, Eq. (1) may be partitioned as